Can you assist me in solving problem...instructions...multiply...5 1/3 * 1 1/4.
Answer by checkley75(3666)

THESE PROBLEMS ARE BEST SOLVED BY CONVERTING THE MIXED FRACTIONS TO IMPROPER FRACTION THUS:
5 1/3=16/3
1 1/4=5/4
NOW MULTIPLY THESE FRACTIONS THUS:
16/3*5/4
16*5/3*4
80/12
6 2/3 OR 20/3 ANSWER.